Find the inverse of the encrypting function



F(x) =(5-x) mod 26 , 0<=x<=25



And use it to decrypt the massage '' ZRQUBFNB''

1
Expert's answer
2022-04-07T11:43:50-0400

The inverse function is the same

"F(x) =(5-x)\\
x=5-F(x)"

"x=(5-25)mod26=6=G\\
x=(5-17)mod26=14=O\\
x=(5-16)mod26=15=P\\
x=(5-20)mod26=11=L\\
x=(5-1)mod26=4=E\\
x=(5-5)mod26=0=A\\
x=(5-13)mod26=18=S\\
x=(5-1)mod26=4=E\\"

Answer: Go please